# Joint Constraints on Fuzzy and Warm Dark Matter from Satellite Populations of the Milky Way and Andromeda

### v1

@@ Abstract
We perform a joint analysis of the Milky Way (MW) and Andromeda (M31) satellite populations to constrain the properties of fuzzy dark matter (FDM) and thermal-relic warm dark matter (WDM). We combine MW satellite observations from the Dark Energy Survey (DES) and Pan-STARRS1 (PS1) with M31 satellite data from the Pan-Andromeda Archaeological Survey (PAndAS), and model the corresponding observable satellite populations using the galaxy–halo connection model together with the appropriate selection functions. Uncertainties in the virial masses of the MW and M31 are incorporated through host-mass priors that linearly scale the relevant model parameters, allowing us to infer the full posterior distributions of all parameters. For the FDM case, we obtain $m_{\mathrm{FDM}}>1.75\times 10^{-20}~\mathrm{eV}$ (95% CL) and $m_{\mathrm{FDM}}>1.41\times 10^{-20}~\mathrm{eV}$ (20:1 posterior ratio). For thermal-relic WDM, we find $m_{\mathrm{WDM}}>6.22~\mathrm{keV}$ (95% CL) and $m_{\mathrm{WDM}}>5.75~\mathrm{keV}$ (20:1 posterior ratio). These results represent a moderate improvement over MW-only constraints, and provide the strongest constraints to date on the FDM and WDM derived from satellite galaxy populations in the Local Group.
@@
software: Astropy (astropy:2013; astropy:2018; astropy:2022), ChainConsumer (Hinton2016), emcee (2013PASP..125..306F), healpy (2005ApJ...622..759G; Zonca2019), Matplotlib (Hunter:2007), NumPy (harris2020array), pandas (mckinney-proc-scipy-2010; reback2020pandas), SciPy (2020SciPy-NMeth)
1 Introduction
@@
Dark matter (DM) constitutes the dominant component of matter in the Universe (2020A&A...641A...6P). Its existence is supported by multiple cosmological and astrophysical observations, such as the cosmic microwave background (CMB; 2020A&A...641A...1P) and galaxy rotation curves (1985ApJ...295..305V), yet its true nature remains elusive. The cold dark matter (CDM) paradigm, exemplified by weakly interacting massive particles (WIMPs), has achieved remarkable success in explaining the large-scale structure (LSS) of the Universe (1996PhR...267..195J; 2018MNRAS.475..624N). However, it faces several well-known small-scale challenges: the missing-satellite problem, the cusp–core problem, and the too-big-to-fail problem (2017Galax...5...17D).
@@
To address these issues and motivated by the lack of any experimental detection of CDM particles, several non-CDM models have been proposed or revisited, such as fuzzy dark matter (FDM; 2000PhRvL..85.1158H; 2021ARA&A..59..247H; 2025arXiv250700705E), warm dark matter (WDM; 2019PrPNP.104....1B), and interacting dark matter (IDM; 2001PhLB..518....8B; 2018PhRvL.121h1301G). These non-CDM scenarios reproduce the LSS of the Universe while simultaneously altering the linear matter power spectrum and the abundance of halos and subhalos on small scales.
@@
Dwarf galaxies, as low-mass and DM-dominated systems, serve as powerful probes of small-scale dark matter structure (2019PhRvL.123e1103M; 2021PhRvL.126i1101N; 2025PhRvL.134o1001Z). Following the success of the Sloan Digital Sky Survey (SDSS) in the early 2000s (2005AJ....129.2692W; 2005ApJ...626L..85W), an increasing number of dwarf galaxies, including progressively fainter systems, have been discovered and continue to be found (2009Natur.461...66M; 2016MNRAS.460.1270D; 2016arXiv161205560C; 2018ApJ...868...55M; 2020ApJ...893...47D; 2022ApJ...933..135D; 2023MNRAS.523..876Q; 2025OJAp....8E..89T; 2025arXiv250704618C; 2025arXiv250912313T), further motivating the use of dwarf galaxies to study dark matter models. Non-CDM models can suppress the abundance of halos and subhalos, and since satellite dwarf galaxies reside within subhalos, it is well-motivated to use their observed abundance to probe the underlying subhalo population, thereby placing constraints on non-CDM models (2021PhRvL.126i1101N; 2021JCAP...08..062N; 2022PhRvD.106l3026D; 2025PhRvD.111f3079T; 2025ApJ...986..127N).
@@
For instance, 2021PhRvL.126i1101N and 2025PhRvD.111f3079T use the galaxy–halo connection model introduced in 2020ApJ...893...48N together with observations of Milky Way (MW) satellites from the Dark Energy Survey (DES; 2016MNRAS.460.1270D) and Pan-STARRS1 (PS1; 2016arXiv161205560C) to constrain various non-CDM models. 2025ApJ...986..127N use the same modeling framework and data, incorporating updated transfer functions and subhalo mass functions to constrain FDM, WDM, and IDM. In addition, 2021JCAP...08..062N and 2022PhRvD.106l3026D develop different semi-analytic frameworks and use the predicted total number of MW satellites from 2018MNRAS.479.2853N and 2020ApJ...893...48N to place constraints on WDM.
@@
In this paper, we present the joint constraints on non-CDM models using the satellite populations of the MW and Andromeda Galaxy (M31). Specifically, we adopt and slightly modify the galaxy–halo connection model introduced by 2020ApJ...893...48N and use observations of MW satellites from DES and PS1 together with M31 satellites from the Pan-Andromeda Archaeological Survey (PAndAS; 2009Natur.461...66M; 2018ApJ...868...55M; 2022ApJ...933..135D; 2023ApJ...952...72D) to constrain the properties of FDM and WDM. Using both satellite populations allows us to improve the overall precision of the constraints, compared to using only the MW satellite population (2024ApJ...967...61N).
@@
This paper is organized as follows: in Section 2, we introduce the FDM and WDM models considered in this work; the observational data of satellite galaxies are presented in Section 3; Section 4 describes the modeling framework; in Section 5, we show the constraint results on the FDM and WDM models; in Section 6, we discuss additional results derived from our modeling framework; we provide a summary and outlook in Section 7. Following 2015ApJ...810...21M and 2025ApJ...986..127N, we adopt the cosmological parameters $h=0.7$, $\Omega_{\mathrm{m}}=0.286$, $\Omega_{\mathrm{b}}=0.049$, $\Omega_{\Lambda}=0.714$, $\sigma_{8}=0.82$, and $n_{\mathrm{s}}=0.96$ (2013ApJS..208...19H). This choice has a negligible impact on our results. Halo masses are defined using the virial overdensity criterion of 1998ApJ...495...80B, which corresponds to $\Delta_{\mathrm{vir}}\approx 99\,\rho_{\mathrm{crit}}$ in our cosmology, where $\rho_{\mathrm{crit}}$ is the critical density of the Universe at $z=0$.
@@
2 Dark Matter Models
@@
2.1 Non-cold dark matter models
@@
where $k$ is the cosmological wave number, $P_{\mathrm{CDM}}(k)$ is the CDM linear matter power spectrum, and $P_{\mathrm{non\text{-}CDM}}(k)$ is the linear matter power spectrum of a non-CDM model (2001ApJ...556...93B).
The transfer functions we consider represent the suppression of the linear matter power spectrum in non-CDM models relative to CDM. To quantify the strength of this suppression, we introduce the half-mode wavenumber $k_{\mathrm{hm}}$, defined as the scale at which the power is reduced to one quarter of that in CDM (2012MNRAS.424..684S), which is given by
@@
In the linear theory, the halo mass and comoving wavenumber are related by (2019ApJ...878L..32N)
@@
where $M$ is the peak virial mass (2025ApJ...986..127N). It should be noted that the specific forms of $\mathcal{T}^{2}_{\mathrm{non\text{-}CDM}}(k)$ and $f_{\mathrm{non\text{-}CDM}}(M)$ depend on the particular dark matter model and its underlying properties. Furthermore, the half-mode mass $M_{\mathrm{hm}}$ derived from $\mathcal{T}^{2}_{\mathrm{non\text{-}CDM}}(k)$ provides a useful characteristic scale that facilitates the construction of $f_{\mathrm{non\text{-}CDM}}(M)$.
2.2 Fuzzy dark matter
FDM consists of ultra-light bosons with mass $m_{\mathrm{FDM}}\sim 10^{-22}~\mathrm{eV}$, whose de Broglie wavelength is on the kiloparsec scale, comparable to the size of dwarf galaxies. Due to quantum pressure, the formation of small-scale structures is suppressed (2000PhRvL..85.1158H; 2021ARA&A..59..247H). Following 2025ApJ...986..127N, we define
@@ 2.2 Fuzzy dark matter
We adopt the FDM transfer function (2022PhRvD.105l3529P)
@@
Using the results from 2025ApJ...986..127N, the suppression of the subhalo mass function in FDM relative to CDM can be written as
@@
2.3 Warm dark matter
In this work, we consider the simplest thermal-relic WDM, whose properties are determined solely by the particle mass $m_{\mathrm{WDM}}$, which is typically a few keV. The non-negligible free-streaming effect of WDM can suppress the formation of small-scale structures (2019PrPNP.104....1B).
Following the same procedure as for FDM, we adopt the WDM transfer function from 2005PhRvD..71f3534V, which is given by
@@ 2.3 Warm dark matter
where we take $\nu=1.049$, and $\alpha(m_{\mathrm{WDM}})$ can be expressed as (2023PhRvD.108d3520V)
@@
Similarly, using the results from 2025ApJ...986..127N, the suppression of the subhalo mass function in WDM relative to CDM can also be expressed as
@@
In Figure 1, we show the examples of the transfer functions and subhalo mass function suppressions for the FDM and WDM models with the same $M_{\mathrm{hm}}$. We can see that, for fixed $M_{\mathrm{hm}}$, the WDM exhibits a stronger suppression than the FDM at relatively high subhalo masses, whereas at lower masses its suppression becomes weaker than that of FDM.
3 Observational data
@@
We consider the observational data of satellite populations in both MW and M31 in the analysis. The observed MW satellite population used in this work is identical to that adopted in 2020ApJ...893...48N. It includes 18 satellites discovered in the PS1 DR1 data (2012ApJ...750...99T) and 16 satellites identified in the DES Y3A2 data (2018ApJS..239...18A), yielding a total of 34 MW satellites. The faintest satellite in this sample, Cetus II, has absolute magnitude $M_{V}\simeq 0~\mathrm{mag}$ (2015ApJ...813..109D).
@@
We use the DES and PS1 survey selection functions derived in 2020ApJ...893...47D, which are publicly available as machine-learning classifiers. These classifiers predict the detection probability of a satellite as a function of its absolute magnitude $M_{V}$, heliocentric distance $D$, azimuthally averaged projected half-light radius $r_{1/2}$, and sky position. In addition, geometric constraints are applied to ensure that observable satellites lie within the survey footprints and to exclude regions where satellite detection is hindered by interstellar extinction, bright nearby stars, or bright extragalactic sources. We only consider satellites located within $300~\mathrm{kpc}$ from the center of the MW. In addition, to exclude likely star clusters, we restrict our sample to satellites with $r_{1/2}>10~\mathrm{pc}$ (2020ApJ...893...48N).
@@
The observed M31 satellite population used in this work is drawn from the PAndAS survey data (2023ApJ...952...72D). We adopt the analytic observational selection function developed by 2022ApJ...933..135D. Although this selection function is analytic rather than a machine-learning classifier, it accounts for the dependence on absolute magnitude $M_{V}$, heliocentric distance $D$, azimuthally averaged projected half-light radius $r_{1/2}$, and sky position, with additional geometric constraints applied.
@@
Thus, although the methods for deriving the MW and M31 selection functions differ, both frameworks are designed to accurately capture the detection probability as a function of the key satellite properties. We therefore do not expect this methodological difference to introduce significant systematic bias into the joint inference of our model parameters.
@@
Specifically, our M31 satellite sample consists of the full set of 24 satellites used in 2023ApJ...952...72D. The faintest satellite in our M31 sample is And XXVI, with an absolute magnitude of $M_{V}\simeq-6.0~\mathrm{mag}$ (2022ApJ...938..101S). Although And XXVI has a very low recovery fraction based on the selection function given in 2022ApJ...933..135D, we have verified that including or excluding this system has a negligible impact on our results. Due to the PAndAS survey footprint, we only consider satellites located between $30$ and $300~\mathrm{kpc}$ from the center of M31. Because of the limitations of the selection function, we additionally restrict our analysis to satellites with $r_{1/2}>63~\mathrm{pc}$, although we have confirmed that our results are insensitive to this cut. In Figure 2, we show the half-light radii $r_{1/2}$ as a function of the absolute magnitudes $M_{V}$ of all 58 satellite galaxies from DES, PS1, and PAndAS used in this work.
4 Modeling Framework

### v2

@@ Abstract
We perform a joint analysis of the Milky Way (MW) and Andromeda (M31) satellite populations to constrain the properties of fuzzy dark matter (FDM) and thermal-relic warm dark matter (WDM). We combine MW satellite observations from the Dark Energy Survey (DES) and Pan-STARRS1 (PS1) with M31 satellite data from the Pan-Andromeda Archaeological Survey (PAndAS), and model the corresponding observable satellite populations using the empirical galaxy–halo connection model described in Nadler et al. (2020) together with the appropriate selection functions. Uncertainties in the virial masses of the MW and M31 are incorporated through host-mass priors that linearly scale the relevant model parameters, allowing us to infer the full posterior distributions of all parameters. For the FDM case, we obtain $m_{\mathrm{FDM}}>1.74\times 10^{-20}~\mathrm{eV}$ (95% CL) and $m_{\mathrm{FDM}}>1.42\times 10^{-20}~\mathrm{eV}$ (20:1 posterior ratio). For thermal-relic WDM, we find $m_{\mathrm{WDM}}>6.20~\mathrm{keV}$ (95% CL) and $m_{\mathrm{WDM}}>5.75~\mathrm{keV}$ (20:1 posterior ratio). These results represent a moderate improvement over MW-only constraints, and provide the strongest constraints to date on the FDM and WDM derived from satellite galaxy populations in the Local Group.
@@
software: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), ChainConsumer (S. R. Hinton, 2016), emcee (D. Foreman-Mackey et al., 2013), healpy (K. M. Górski et al., 2005; A. Zonca et al., 2019), Matplotlib (J. D. Hunter, 2007), NumPy (C. R. Harris et al., 2020), pandas (Wes McKinney, 2010; T. pandas development team, 2020), SciPy (P. Virtanen et al., 2020)
I Introduction
@@
Dark matter (DM) constitutes the dominant component of matter in the Universe (Planck Collaboration et al., 2020a). Its existence is supported by multiple cosmological and astrophysical observations, such as the cosmic microwave background (CMB; Planck Collaboration et al. 2020b) and galaxy rotation curves (V. C. Rubin & W. K. Ford, 1970; T. S. van Albada et al., 1985; Y. Sofue, 2015), yet its true nature remains elusive. The cold dark matter (CDM) paradigm, exemplified by weakly interacting massive particles (WIMPs), has achieved remarkable success in explaining the large-scale structure (LSS) of the Universe (G. Jungman et al., 1996; V. Springel et al., 2005; C. S. Frenk & S. D. M. White, 2012; M. Vogelsberger et al., 2014; D. Nelson et al., 2018). However, it faces several well-known small-scale challenges: the missing-satellite problem, the cusp–core problem, and the too-big-to-fail problem (A. Del Popolo & M. Le Delliou, 2017; J. S. Bullock & M. Boylan-Kolchin, 2017; S. Vegetti et al., 2026).
@@
To address these issues and motivated by the lack of any experimental detection of CDM particles, several non-CDM models have been proposed or revisited, such as fuzzy dark matter (FDM; W. Hu et al. 2000; L. Hui 2021; A. Eberhardt & E. G. M. Ferreira 2025), warm dark matter (WDM; A. Boyarsky et al. 2019), and interacting dark matter (IDM; C. Bœhm et al. 2001; V. Gluscevic & K. K. Boddy 2018). These non-CDM scenarios reproduce the LSS of the Universe while simultaneously altering the linear matter power spectrum and the abundance of halos and subhalos on small scales. We should note that the role of baryonic physics is also significant at small scales, which can alleviate some of the small-scale challenges faced by the CDM model without the necessity of invoking non-CDM (A. Di Cintio et al., 2014; T. K. Chan et al., 2015; J. S. Bullock & M. Boylan-Kolchin, 2017; L. V. Sales et al., 2022). Therefore, the main reason for these issues remains a subject of ongoing debate and is yet to be fully resolved (J. S. Bullock & M. Boylan-Kolchin, 2017; L. V. Sales et al., 2022).
@@
Dwarf galaxies, as low-mass and DM-dominated systems, serve as powerful probes of small-scale dark matter structure (D. J. E. Marsh & J. C. Niemeyer, 2019; E. O. Nadler et al., 2021b; T. Zimmermann et al., 2025). Following the success of the Sloan Digital Sky Survey (SDSS) in the early 2000s (B. Willman et al., 2005a, b), an increasing number of dwarf galaxies, including progressively fainter systems, have been discovered and continue to be found (A. W. McConnachie et al., 2009; Dark Energy Survey Collaboration et al., 2016; K. C. Chambers et al., 2016; A. W. McConnachie et al., 2018; A. Drlica-Wagner et al., 2020; A. Doliva-Dolinsky et al., 2022; H. Qu et al., 2023; K. Tsiane et al., 2025; CSST Collaboration et al., 2025; C. Y. Tan et al., 2025b), further motivating the use of dwarf galaxies to study dark matter models. Non-CDM models can suppress the abundance of halos and subhalos, and since satellite dwarf galaxies reside within subhalos, it is well-motivated to use their observed abundance to probe the underlying subhalo population, thereby placing constraints on non-CDM models (E. O. Nadler et al., 2021b; O. Newton et al., 2021; A. Dekker et al., 2022; C. Y. Tan et al., 2025a; E. O. Nadler et al., 2025).
@@
For instance, E. O. Nadler et al. (2021b) and C. Y. Tan et al. (2025a) use the empirical galaxy–halo connection model introduced in E. O. Nadler et al. (2020) together with observations of Milky Way (MW) satellites from the Dark Energy Survey (DES; Dark Energy Survey Collaboration et al. 2016) and Pan-STARRS1 (PS1; K. C. Chambers et al. 2016) to constrain various non-CDM models. E. O. Nadler et al. (2025) use the same modeling framework and data, incorporating updated transfer functions and subhalo mass functions to constrain FDM, WDM, and IDM. In addition, O. Newton et al. (2021) and A. Dekker et al. (2022) develop different semi-analytic frameworks and use the predicted total number of MW satellites from O. Newton et al. (2018) and E. O. Nadler et al. (2020) to place constraints on WDM.
@@
In this paper, we present the joint constraints on non-CDM models using the satellite populations of the MW and Andromeda Galaxy (M31). Specifically, we adopt and slightly modify the empirical galaxy–halo connection model introduced by E. O. Nadler et al. (2020) and use observations of MW satellites from DES and PS1 together with M31 satellites from the Pan-Andromeda Archaeological Survey (PAndAS; A. W. McConnachie et al. 2009, 2018; A. Doliva-Dolinsky et al. 2022, 2023) to constrain the properties of FDM and thermal-relic WDM. Using both satellite populations allows us to improve the overall precision of the constraints, compared to using only the MW satellite population (E. O. Nadler et al., 2024).
@@
This paper is organized as follows: in Section II, we introduce the FDM and WDM models considered in this work; the observational data of satellite galaxies are presented in Section III; Section IV describes the modeling framework; in Section V, we show the constraint results on the FDM and WDM models; in Section VI, we discuss additional results derived from our modeling framework; we provide a summary and outlook in Section VII. Following Y.-Y. Mao et al. (2015) and E. O. Nadler et al. (2025), we adopt the cosmological parameters $h=0.7$, $\Omega_{\mathrm{m}}=0.286$, $\Omega_{\mathrm{b}}=0.049$, $\Omega_{\Lambda}=0.714$, $\sigma_{8}=0.82$, and $n_{\mathrm{s}}=0.96$ (G. Hinshaw et al., 2013). This choice has a negligible impact on our results. Halo masses are defined using the virial overdensity criterion of G. L. Bryan & M. L. Norman (1998), which corresponds to $\Delta_{\mathrm{vir}}\approx 99\,\rho_{\mathrm{crit}}$ in our cosmology, where $\rho_{\mathrm{crit}}$ is the critical density of the Universe at $z=0$.
@@
II Dark Matter Models
@@
II.1 Non-cold dark matter models
@@
where $k$ is the cosmological wave number, $P_{\mathrm{CDM}}(k)$ is the CDM linear matter power spectrum, and $P_{\mathrm{non\text{-}CDM}}(k)$ is the linear matter power spectrum of a non-CDM model (P. Bode et al., 2001).
The transfer functions we consider represent the suppression of the linear matter power spectrum in non-CDM models relative to CDM. To quantify the strength of this suppression, we introduce the half-mode wavenumber $k_{\mathrm{hm}}$, defined as the scale at which the power is reduced to one quarter of that in CDM (P. Bode et al., 2001; A. Schneider et al., 2012), which is given by
@@
In the linear theory, the halo mass and comoving wavenumber are related by (P. Bode et al., 2001; A. Schneider et al., 2012; E. O. Nadler et al., 2019a)
@@
where $M$ is the peak virial mass (E. O. Nadler et al., 2025). It should be noted that the specific forms of $\mathcal{T}^{2}_{\mathrm{non\text{-}CDM}}(k)$ and $f_{\mathrm{non\text{-}CDM}}(M)$ depend on the particular dark matter model and its underlying properties. Furthermore, for a specific non-CDM model, the half-mode mass $M_{\mathrm{hm}}$ derived from $\mathcal{T}^{2}_{\mathrm{non\text{-}CDM}}(k)$ provides a useful characteristic scale that facilitates the construction of $f_{\mathrm{non\text{-}CDM}}(M)$. However, using $M_{\mathrm{hm}}$ for a direct comparison among different non-CDM models can be an oversimplification and may introduce significant biases in the resulting constraints (R. Murgia et al., 2017, 2018; D. C. Hooper et al., 2022).
II.2 Fuzzy dark matter
FDM consists of ultra-light bosons with mass $m_{\mathrm{FDM}}\sim 10^{-22}~\mathrm{eV}$, whose de Broglie wavelength is on the kiloparsec scale, comparable to the size of dwarf galaxies. Due to quantum pressure, the formation of small-scale structures is suppressed (W. Hu et al., 2000; L. Hui, 2021). Following E. O. Nadler et al. (2025), we define
@@ II.2 Fuzzy dark matter
We adopt the FDM transfer function (S. Passaglia & W. Hu, 2022)
@@
Using the results from E. O. Nadler et al. (2025), the suppression of the subhalo mass function in FDM relative to CDM can be written as
@@
II.3 Warm dark matter
In this work, we consider the simplest thermal-relic WDM, whose properties are determined solely by the particle mass $m_{\mathrm{WDM}}$, which is typically a few keV. The non-negligible free-streaming effect of WDM can suppress the formation of small-scale structures (A. Boyarsky et al., 2019). We should notice that, although thermal-relic WDM is frequently employed as a benchmark for non-CDM models, the functional forms of the transfer functions for other scenarios may differ from the thermal-relic case (see e.g. I. A. Zelko et al., 2022).
Following the same procedure as for FDM, we adopt the WDM transfer function from M. Viel et al. (2005), which is given by
@@ II.3 Warm dark matter
where we take $\nu=1.049$, and $\alpha(m_{\mathrm{WDM}})$ can be expressed as (C. M. Vogel & K. N. Abazajian, 2023)
@@
Similarly, using the results from E. O. Nadler et al. (2025), the suppression of the subhalo mass function in WDM relative to CDM can also be expressed as
@@
It is worth noting that when characterizing the suppression of the subhalo mass function in FDM and WDM relative to CDM, E. O. Nadler et al. (2025) utilized cosmological zoom-in simulations with modified initial conditions corresponding to FDM or WDM transfer functions. Although these simulations neglect FDM wave interference effects and the thermal velocities of WDM particles, this approximation is sufficiently accurate for the subhalos of interest in this work, which have peak virial masses of $M\gtrsim 10^{8}~M_{\odot}$ (M. Leo et al., 2017; E. O. Nadler et al., 2020; S. May & V. Springel, 2023). The probabilistic inference was performed to reconstruct the model parameters ($\alpha$, $\beta$, and $\gamma$) which fit the simulated data, and the values that maximize the posterior distributions were adopted. They verified that the uncertainties associated with these parameters do not significantly affect the final constraints. In Figure 1, we show the examples of the transfer functions and subhalo mass function suppressions for the FDM and WDM models with the same $M_{\mathrm{hm}}$. We can see that, for fixed $M_{\mathrm{hm}}$, the WDM exhibits a stronger suppression than the FDM at relatively high subhalo masses, whereas at lower masses its suppression becomes weaker than that of FDM.
@@ III Observational data
We consider the observational data of satellite populations in both MW and M31 in the analysis. The observed MW satellite population used in this work is identical to that adopted in E. O. Nadler et al. (2020). It includes 18 satellites discovered in the PS1 DR1 data (J. L. Tonry et al., 2012) and 16 satellites identified in the DES Y3A2 data (T. M. C. Abbott et al., 2018), yielding a total of 34 MW satellites. The faintest satellite in this sample, Cetus II, has absolute magnitude $M_{V}\simeq 0~\mathrm{mag}$ (A. Drlica-Wagner et al., 2015).
@@
We use the DES and PS1 survey selection functions derived in A. Drlica-Wagner et al. (2020), which are publicly available as machine-learning classifiers. [^1] https://github.com/des-science/mw-sats These classifiers predict the detection probability of a satellite as a function of its absolute magnitude $M_{V}$, heliocentric distance $D$, azimuthally averaged projected half-light radius $r_{1/2}$, and sky position. In addition, geometric constraints are applied to ensure that observable satellites lie within the survey footprints and to exclude regions where satellite detection is hindered by interstellar extinction, bright nearby stars, or bright extragalactic sources. We only consider satellites located within $300~\mathrm{kpc}$ from the center of the MW. In addition, to exclude likely star clusters, we restrict our sample to satellites with $r_{1/2}>10~\mathrm{pc}$ (E. O. Nadler et al., 2020).
@@
The observed M31 satellite population used in this work is drawn from the PAndAS survey data (A. Doliva-Dolinsky et al., 2023). We adopt the analytic observational selection function developed by A. Doliva-Dolinsky et al. (2022). [^2] https://github.com/dolivadolinsky/Recovery_dwarf_galaxy_M31 Although this selection function is analytic rather than a machine-learning classifier, it accounts for the dependence on absolute magnitude $M_{V}$, heliocentric distance $D$, azimuthally averaged projected half-light radius $r_{1/2}$, and sky position, with additional geometric constraints applied.
@@
Thus, although the methods for deriving the MW and M31 selection functions differ, both frameworks are designed to accurately capture the detection probability as a function of the key satellite properties. Moreover, the joint constraints are not dominated by the M31 selection function. We therefore do not expect this methodological difference to introduce significant systematic bias into the joint inference of our model parameters. We have also verified through sensitivity test that the impact of this methodological difference is minimal.
@@
Specifically, our M31 satellite sample consists of the full set of 24 satellites used in A. Doliva-Dolinsky et al. (2023). The faintest satellite in our M31 sample is And XXVI, with an absolute magnitude of $M_{V}\simeq-6.0~\mathrm{mag}$ (A. Savino et al., 2022). Although And XXVI has a very low recovery fraction based on the selection function given in A. Doliva-Dolinsky et al. (2022), we have verified that including or excluding this system has a negligible impact on our results. Due to the PAndAS survey footprint, we only consider satellites located between $30$ and $300~\mathrm{kpc}$ from the center of M31. Because of the limitations of the selection function, we additionally restrict our analysis to satellites with $r_{1/2}>63~\mathrm{pc}$, although we have confirmed that our results are insensitive to this cut. In Figure 2, we show the half-light radii $r_{1/2}$ as a function of the absolute magnitudes $M_{V}$ of all 58 satellite galaxies from DES, PS1, and PAndAS used in this work.
IV Modeling Framework